Back to the NIKA2 Pool page

Daily Reports

March 19th (Tuesday)

Bilal, Alessia, Stefano, Juan, Isabelle, Pablo T. arrived.

[Afternoon shift: Bilal & Stefano]

18h00 UT started with V3 acquisition after maintenance delayed because of VAC1 problems. Unstable and cloudy sky.


[Night shift: Juan & Pablo]

Very good weather conditions, opacity around 0.1. We performed observations for LPSZ , GASTON and Saturn/Iapetus

March 20th (Wednesday)

[Morning shift: Isabelle & Alessia]

Observing with good opacity and clear sky until ~11h20. Observed project 111-18, 159-18. Bad sky, unstable, cloudy since then. Telescope left to Juan P. to perform technical tests.

[Afternoon shift: Bilal & Stefano]



[Night shift: Juan & Pablo] The wind was too strong at the beginning of the night and we could not start observations until 3h15 UTC.

March 21st (Thursday)

Samuel arrived, Pablo left.

[Morning shift: Isabelle & Alessia]

Observing in very good conditions until 11h UTC. Projects observed: GASTON, 111-18 and 159-18. Observations disturbed by the ski station snow gun. Pictures below.



[Afternoon shift: Bilal & Stefano]

Technical tests (nikaw-18) while beams not great due to sky conditions. Then 192-16 (GOODSNORTH).


[Night shift: Juan & Samuel]

Observations continues with DAQV3 old parameters. Mostly 199-16 (GT_LP SZ) and 122-16 (GT-LP GASTON), plus some additional nikaw18 for completing V3 checklist.

March 22nd (Friday)

[Morning shift: Alessia & Isabelle]

6h40 UTC: change to DAQV1 with new parameters. From scan #75 on.

Several back and forth (morning and afternoon) between old and new param files because finding the new geometry and implementing it correctly is not straightforward; it's necessary to have a good beammaps, to remove bad pixels on the new geometry, to find correctly new pixel names and positions, find reference pixels, match arrays old and new geometries, check pointing, debug possible offsets, and so on...

[Afternoon shift: Bilal & Stefano]

We switch to new parameter file for A1 and A3. Did 3 beammap and switch back to the old parameter until full treatment of the beammaps.

Meanwhile Juan - with some help by Stefano - solves two long-standing issues: now v3 imbfits contain the sweep frequencies and the signal is positive.


[Night shift: Juan & Samuel]

Weather forecast not great for the coming days, few short nice slots ahead of us and then big degradation. As we didn't totally solved the new geometry we decide to stay on the safe side of the old parameters to be sure to not spoil the nice weather slots and move forward with the pool. Weather conditions were unstable at the beginning of the night, so we mainly concentrated in project 193-18 which is not very demanding with sky conditions. Then we switched to 199-16 for 1h, then to 184-18.

March 23rd (Saturday)

[Morning shift: Alessia & Isabelle]

Good weather conditions until 10h40 UTC, 159-18 project observed. We have tried to switch to EMIR for a 3mm project but it was not possible to observe. Restart of the observations at 16h30 UTC.

[Afternoon shift: Bilal & Stefano]

148-18 for 1.5 hours.

[Night shift: Juan & Samuel]

207-18 for ~1h in weather condition starting to degrades due to increasing wind. Had to stop and park the antenna at 23h UT dut to wind too strong. Then clouds arrived and weather got bad.

March 24th (Sunday)

[Morning shift: Alessia & Isabelle]


[Afternoon shift: Bilal & Stefano]

Snow and then clouds

It clears up in the evening and we start observing at 20:56 UT, we pass the baton to the next shift.

[Night shift: Juan & Samuel]

We attempt several times to perform pointing and focus. We also try some calibration on Pallas. Then we observe 193-18 and 184-18 projects.

March 25th (Monday)

Pablo arrived

[Morning shift: Alessia & Isabelle]

[Afternoon shift: Bilal & Stefano]

A nice sunset for our season's last evening's goodbye:


DailyReportsNika2Pool190319 (last edited 2019-03-25 21:02:26 by NikaBolometer)